A Bibliography of Supercomputing ’88

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.02

Abstract

This bibliography records articles presented at the Supercomputing ’88 conference.

Title word cross-reference

95800^{4} + 217519^{4} + 414560^{4} = 422481^{4} [Fry88].

1 [HSS88]. 1-port [hTDT88]. 10 [BK88, DSS88].

2 [BTR88, EJL88, SS88, SW88, Swi88, TSM88, LMT88]. 200 [BK88]. 205 [BK88].

3D [PS88].

416 [SW88].

80 [EML88]. 820/80 [EML88]. ’88 [IEE88, ML88].

Adaptation [RS88]. aerodynamics [PB88]. algorithm [AG88, BF88, Bis88, NPD88, RS88, Swi88]. algorithms [Dub88, OG88, Yos88]. alignments [SD88b]. Analysis [PS88, Aki88, IMS88]. ANSYS [HMB^{+}88]. applications [CC88, GGGJ88, Hol88]. approach [Yos88]. architecture [CGMW88, DG88, KS88, PS88, TS88]. area [WL88]. ARGOT [OS88]. arithmetic [Kan88]. array [LK88]. arrays [DC88, GT88]. Assessing [ACM88]. assessment [CF88]. associative [SWC88]. automata [BTR88]. automatic [TK88].

based [RKZ88]. BBN [OS88]. Benchmarking [GJ88, SS88]. benefits [ACM88]. birth [EML88]. block [Dub88, EJL88]. Boeing [GKS88]. bottlenecks [IL88]. Breakthroughs [SC88]. buffer [FM88]. Butterfly [OS88].

C [SW88]. C-MP [SW88]. C-MP/416 [SW88]. C240 [CGMW88]. cache [SD88a]. calculation [Kan88]. Carlo [PB88]. Cedar [GPHL88]. cellular [BTR88]. census [Bur88]. CFD [Aki88]. chaining [hTDT88]. change [CF88, Rob88, Wor88]. change-human [CF88]. chemistry [WP88]. climate [SC88]. clustering [RS88]. CM [BTR88, LMT88]. CM-2 [BTR88, LMT88]. coalitions [SD88b]. code [IMS88]. comparison [LMT88, SW88]. compilation [BK88]. compiler [TK88]. compilers [CDL88]. Compiling [Dra88, GP88, TIGU88]. Complementary [Bur88]. compression [BRM88]. computational [BBM88, Hol88]. computations [CC88]. Computer [WP88, Abe88, BTR88, Bur88, HMB^{+}88, TIGU88]. computers [GJ88, Wor88]. computing [BCC^{+}88, LK88]. Congress [SD88b]. conjugate [NPR88]. conjugate-gradient [NPR88]. Connection [BTR88, Fry88, LMT88, SS88, TSM88]. Convex [CGMW88]. cosmic [Mat88]. Council [DEDR88]. CPU [EJL88]. Cray [EJL88, FY88, SW88, Swi88]. Cray-2 [EJL88, SW88, Swi88]. CRegs [DC88]. Cyber [BK88].

data [BRM88, SWC88]. database [Bur88]. dataflow [ACM88, DG88, HSS88, IL88]. decimal [Kan88]. defined [GT88]. Delaunay [FY88]. derivation [Yos88]. Design [FM88, Nev88, SD88a]. Development [HTSH88, AG88]. dialects [GPHL88]. digits [Kan88]. dimensional [FY88, OG88, Swi88]. discord [Nev88]. disk [IH88]. display [SSG88]. Distributed [JHH^{+}88]. dynamic [IL88]. dynamics [Hol88].

Efficient [HSS88, DG88]. efforts [Aki88]. Elimination [IL88]. engine [WP88]. environment [GGGJ88]. equations [TSM88, Yos88]. ETA [BK88, DSS88]. ETA-10 [BK88]. evaluation [BK88]. execution [RS88]. experiments [BBM88]. extensions [TK88].

FACOM [IH88]. factorization [Bis88]. fast [CC88]. Faust [GGGJ88]. Finding [Fry88]. fine [ACM88]. fine-grain [ACM88]. first [TIGU88]. first-order [TIGU88]. Florida [IEE88, ML88]. flow [Aki88, BTR88, SWC88]. fluid [Hol88]. form [Dub88]. Fortran [BK88, GPHL88]. Fourier [CC88]. frame [FM88]. framework [CF88]. free [SD88a].

generated [BK88]. generation [EML88]. Global [CF88, Rob88]. gradient [NPR88]. grain [ACM88]. graphics [Mat88]. Growing [Nev88].

hardware [Nev88, Yos88]. Hessenberg [Dub88]. high [BF88, BCC^{+}88, LK88, SD88a, Wor88]. high-performance [SD88a, Wor88]. high-speed [BF88, BCC^{+}88, LK88]. Highly [NPD88]. Hitachi [EML88]. Horizon [CC88, Dra88, Gle88, Kop88, KS88, TS88]. HORSE [Kop88]. Householder [Dub88]. human [CF88]. hyperplane [GT88].

I-NET [WW88]. III [OS88]. impact [GJ88]. implementation [FM88]. incompressible [TSM88]. influence [Rob88]. instructions [WW88]. integrated [BCC^{+}88]. integration [GKS88]. interaction [Mat88]. interactions [CF88]. Interactive [SSG88]. inversion [BF88]. IPP [Abe88]. Ising [DSS88]. ISODATA [RS88]. issue [SD88b]. issues [GP88]. issuing [WW88]. iWarp [BCC^{+}88]. job [HTSH88]. job-job [HTSH88]. Joint [DEDR88].

kernel [HMB^{+}88]. kind [DC88]. knock [WP88].

language [TK88]. large [NPR88]. large-scale [NPR88]. Linda [WL88]. linear [LK88, TIGU88]. local [Bis88, WL88]. lockup [SD88a]. lockup-free [SD88a]. LSOLVE [HMB^{+}88].

Machine [SS88, TSM88, BTR88, Fry88, LMT88]. made [SC88]. magneto [IH88]. magneto-optic [IH88]. making [JHH^{+}88]. Mass [Bed88, CDK88, RKZ88]. matrix [BF88, CC88, Dub88]. mechanism [WW88]. memory [DC88, PS88]. methods [NPR88]. metrics [LMT88]. minimization [NPR88]. mixed [LJ88]. model [DSS88]. modeling [SC88, WP88]. molecular [PB88]. Monte [PB88]. movie [JHH^{+}88]. MP [FY88]. MP/416 [SW88]. Multi [SD88b]. Multi-issue [SD88b]. multiple [Kan88, WW88]. multiple-precision [Kan88]. multiprocessor [LJ88, OS88]. multiprocessors [SD88a]. Multitasking [DSS88, Swi88]. multivariate [CF88].

NAL [Aki88, HTSH88]. Navier [Swi88, TSM88]. NET [WW88]. network [PS88, WL88]. Newton [BF88]. no [TK88]. nonstationary [Aki88]. November [IEE88, ML88]. nuclear [IMS88]. numerical [HTSH88].

ocean [SC88]. optic [IH88]. Optimal [LJ88]. order [TIGU88]. Orlando [IEE88, ML88]. other [GPHL88].

Parallel [AG88, BBM88, Bis88, BCC^{+}88, GJ88, GGGJ88, GPHL88, HMB^{+}88, IMS88, SSG88]. parallelism [ACM88]. parallelizable [BF88]. Pascal [TK88]. patterns [Wor88]. Performance [Gle88, SW88, EJL88, GJ88, SD88a, Wor88]. personal [Bur88]. phase [BTR88]. philosophy [Nev88]. physical [GJ88]. pi [Kan88]. pipelined [DG88]. pivoting [Bis88]. plumes [OG88]. pointers [DC88]. policies [LJ88]. Polycyclic [hTDT88]. port [hTDT88]. possible [SC88]. pre [BK88]. pre-compilation [BK88]. precision [Kan88]. prediction [Gle88]. Present [Abe88]. problem [PB88]. Proceedings [IEE88, ML88]. processing [BBM88, Bur88, HSS88, IMS88]. processor [DG88, TS88]. processors [IL88]. Profiles [CDK88]. program [GKS88, HMS88, Kan88]. programmable [LK88]. programming [GGGJ88, Nev88]. programs [ACM88]. project [Bur88]. protein [LMT88].

QR [Bis88].

Rayleigh [PB88]. reactor [IMS88, NPD88]. recurrence [Yos88]. recurrences [TIGU88]. recursively [GT88]. reduction [Dub88]. referencing [DC88]. RELAP5 [IMS88]. Research [DEDR88, Mat88]. results [CDL88].

S [EML88]. S-820 [EML88]. S-820/80 [EML88]. scalar [LJ88]. scalar-vector [LJ88]. scale [NPR88]. scheduler [HTSH88]. scheduling [EJL88, LJ88, hTDT88]. scheme [BRM88]. scientific [GGGJ88, HMS88, JHH^{+}88, Phi88, SSG88]. second [EML88]. sequence [LMT88]. set [BBM88]. shared [PS88]. SIGMA [HSS88]. SIGMA-1 [HSS88]. simulation [Abe88, BTR88, GJ88, Kop88, NPD88, OG88, PB88]. simulator [HTSH88]. software [GJ88, KS88]. solar [Rob88]. solution [BCC^{+}88, TSM88]. Some [Wor88]. space [NPD88]. sparse [CC88]. Spectral [TSM88]. speed [BF88, BCC^{+}88, LK88]. Squeezing [EJL88]. static [SWC88]. status [Abe88]. step [HTSH88]. Stokes [Swi88, TSM88]. storage [Bed88, CDK88, RKZ88]. Strassen [BF88]. Strassen-Newton [BF88]. strategy [OS88]. strings [Mat88]. Study [LMT88, CC88, Mat88]. sub [IH88]. sub-system [IH88]. suite [CDL88]. Supercomputer [GKS88, Hol88, Bur88, CC88, DSS88, FM88, Gle88, HSS88, Kop88, RS88, SD88b]. supercomputer/personal [Bur88]. supercomputers [BBM88, BRM88, GP88, LJ88, RKZ88, SC88, SW88, hTDT88]. Supercomputing [Aki88, IEE88, ML88, Bed88, DEDR88, HMS88, KS88, WL88]. support [Bed88]. symbolic [GT88]. system [FM88, IH88, KS88, RKZ88]. systems [CDK88, NPD88].

tale [CDK88]. teaching [Mat88]. techniques [SSG88, TIGU88]. technological [Wor88]. templates [SWC88]. test [CDL88]. Three [FY88, OG88, Swi88]. Three-dimensional [OG88, Swi88]. today [SC88]. Tokenless [SWC88]. tomorrow [SC88]. tools [BK88]. toroidal [PS88]. training [HMS88]. transform [CC88]. transformation [GT88]. transformational [Yos88]. transient [IMS88, NPD88]. triangulation [FY88]. Tuning [HMB^{+}88]. two [BTR88, CC88, CDK88]. two-phase [BTR88].

underexpanded [OG88]. unit [DEDR88]. Unix [RKZ88]. upwind [OG88]. users [HMS88]. Using [WL88, Bis88, OG88, SWC88].

V [TK88]. V-Pascal [TK88]. Vector [IMS88, BK88, EJL88, GJ88, GPHL88, HSS88, LJ88, RS88, TIGU88, hTDT88]. Vectorization [Kan88, NPR88]. Vectorized [PB88, NPD88]. Vectorizing [CDL88, TK88]. video [JHH^{+}88]. visualization [Phi88, SSG88]. voting [SD88b]. vs [hTDT88].

workbench [AG88, Phi88].

X [FY88]. X-MP [FY88].

References

Abe:1988:PSC

[Abe88] Y. Abe. Present status of computer simulation at IPP. In Martin and Lundstrom [ML88], pages 72–80. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Arvind:1988:ABF

[ACM88] Arvind, D. E. Culler, and G. K. Maa. Assessing the benefits of fine-grain parallelism in dataflow programs. In IEEE [IEE88], pages 60–69. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Arya:1988:PAD

[AG88] S. Arya and B. Gaither. Parallel algorithm development workbench. In IEEE [IEE88], pages 11–17. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Aki:1988:SEC

[Aki88] T. Aki. Supercomputing efforts on CFD for nonstationary flow analysis at NAL. In Martin and Lundstrom [ML88], pages 61–70. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Balram:1988:PPS

[BBM88] N. Balram, C. Belo, and J. M. F.

Moura. Parallel processing on supercomputers: a set of computational experiments. In IEEE [IEE88], pages 247–257. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Borkar:1988:IIS

[BCC^{+}88] S. Borkar, R. Cohn, G. Cox, S. Gleason, T. Gross, Kung, H. T., M. Lam, B. Moore, C. Peterson, J. Pieper, Rankin, L., P. S. Tseng, J. Sutton, J. Urbanski, and J. Webb. iWarp: an integrated solution to high-speed parallel computing. In IEEE [IEE88], pages 330–339. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Bedoll:1988:MSS

[Bed88] R. F. Bedoll. Mass storage support for supercomputing. In IEEE [IEE88], pages 217–221. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Bailey:1988:SAH

[BF88] D. H. Bailey and H. R. P. Ferguson. A Strassen-Newton algorithm for high-speed parallelizable matrix inversion. In IEEE [IEE88], pages 419–424. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Bischof:1988:PQF

[Bis88] C. H. Bischof. A parallel QR factorization algorithm using local pivoting. In IEEE [IEE88], pages 400–407. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Braswell:1988:EVF

[BK88] R. N. Braswell and M. S. Keech.

An evaluation of vector Fortran 200 generated by Cyber 205 and ETA-10 pre-compilation tools. In IEEE [IEE88], pages 106–113. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Bassiouni:1988:SDC

[BRM88] M. A. Bassiouni, N. Ranganathan, and A. Mukherjee. A scheme for data compression in supercomputers. In IEEE [IEE88], pages 272–278. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Boghosian:1988:CAS

[BTR88] B. M. Boghosian, W. Taylor, IV, and D. H. Rothman. A cellular automata simulation of twophase flow on the CM-2 Connection Machine computer. In Martin and Lundstrom [ML88], pages 34–44. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Burton:1988:CPS

[Bur88] O. V. Burton. Complementary processing: a supercomputer/personal computer US census database project. In Martin and Lundstrom [ML88], pages 167–177. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Carlson:1988:FFT

[CC88] D. A. Carlson and J. M. Conroy. The fast Fourier transform and sparse matrix computations: a study of two applications on the Horizon supercomputer. In IEEE [IEE88], pages 55–58. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Collins:1988:PMS

[CDK88] B. Collins, M. Devaney, and D. Kitts. Profiles in mass storage: a tale of two systems. In IEEE [IEE88], pages 222–228. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Callahan:1988:VCT

[CDL88] David Callahan, Jack Dongarra, and David Levine. Vectorizing compilers: A test suite and results. In IEEE [IEE88], pages 98–105. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Chen:1988:GCI

[CF88] R. S. Chen and M. B. Fiering. Global change-human interactions: a multivariate assessment framework. In Martin and Lundstrom [ML88], pages 251–261. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Chastain:1988:CCA

[CGMW88] M. Chastain, G. Gostin, J. Mankovich, and S. Wallach. The Convex C240 architecture. In IEEE [IEE88], pages 321–329. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Dietz:1988:CNK

[DC88] H. Dietz and C.-H. Chi. CRegs: a new kind of memory for referencing arrays and pointers. In IEEE [IEE88], pages 360–367. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Davies:1988:JRC

[DEDR88] B. W. Davies, R. G. Evans, T. Daniels, and D. J. Rigby. The Joint Research Council’s supercomputing unit. In IEEE [IEE88], pages 450–455. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v.

2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Dennis:1988:EPD

[DG88] J. B. Dennis and G. R. Gao. An efficient pipelined dataflow processor architecture. In IEEE [IEE88], pages 368–373. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Draper:1988:CH

[Dra88] J. M. Draper. Compiling on Horizon. In IEEE [IEE88], pages 51–52. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Duke:1988:MIM

[DSS88] D. W. Duke, R. Salvador, and D. Sandee. Multitasking the Ising model on the ETA 10 supercomputer. In Martin and Lundstrom [ML88], pages 84–94. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Dubrulle:1988:BHA

[Dub88] A. A. Dubrulle. On block Householder algorithms for the reduction of a matrix to Hessenberg form. In Martin and Lundstrom [ML88], pages 129–140. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Eisenbeis:1988:SMC

[EJL88] C. Eisenbeis, W. Jalby, and A. Lichnewsky. Squeezing more CPU performance out of a Cray-2 by vector block scheduling. In IEEE [IEE88], pages 237– 245. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Eoyang:1988:BSG

[EML88] C. Eoyang, R. H. Mendez, and O. M. Lubeck. The birth of the second generation: the Hitachi S-820/80. In IEEE [IEE88], pages 296–303. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Fowler:1988:DIS

[FM88] J. D. Fowler, Jr. and M. Mc-Gowen. Design and implementation of a supercomputer frame buffer system. In IEEE [IEE88], pages 140–147. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Frye:1988:FCM

[Fry88] R. E. Frye. Finding 95800^{4} + 217519^{4} + 414560^{4} = 422481^{4} on the Connection Machine. In Martin and Lundstrom [ML88], pages 106–116. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Field:1988:TDD

[FY88] D. A. Field and K. Yarnall.

Three dimensional Delaunay triangulation on a Cray X-MP. In Martin and Lundstrom [ML88], pages 202–213. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Guarna:1988:FEP

[GGGJ88] V. A. Guarna, Jr., D. Gannon, Y. Gaur, and D. Jablonowski. Faust: an environment for programming parallel scientific applications. In IEEE [IEE88], pages 3–10. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Ginsberg:1988:BPP

[GJ88] M. Ginsberg and J. P. Johnson. Benchmarking the performance of physical impact simulation software on vector and parallel computers. In Martin and Lundstrom [ML88], pages 180–190. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Griffin:1988:SIP

[GKS88] R. B. Griffin, J. S. Kowalik, and M. R. Scott. Supercomputer integration program at Boeing. In IEEE [IEE88], pages 442– 449. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Glenn:1988:PPH

[Gle88] R. R. Glenn. Performance prediction for the Horizon supercomputer. In IEEE [IEE88], pages 48–50. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Girkar:1988:CIS

[GP88] M. Girkar and C. Polychronoupoulos. Compiling issues for supercomputers. In IEEE [IEE88], pages 164–173. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Guzzi:1988:CFO

[GPHL88] M. D. Guzzi, D. A. Padua, J. P.

Hoeflinger, and D. H. Lawrie. Cedar Fortran and other vector and parallel Fortran dialects. In IEEE [IEE88], pages 114– 121. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Gokhale:1988:SHT

[GT88] M. B. Gokhale and T. C. Torgersen. The symbolic hyperplane transformation for recursively defined arrays. In IEEE [IEE88], pages 207–214. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Hessel:1988:TAK

[HMB^{+}88] R. E. Hessel, M. Myszewski, G. Brussino, J. A. Swanson, and L. Wagner. Tuning the ANSYS kernel LSOLVE for a parallel computer. In Martin and Lundstrom [ML88], pages 191–201. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Hanson:1988:TPS

[HMS88] F. Hanson, T. Moher, and N. Sabelli. A training program for scientific supercomputing users. In IEEE [IEE88], pages 342– 349. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Holst:1988:SAC

[Hol88] T. L. Holst. Supercomputer applications in computational fluid dynamics. In Martin and Lundstrom [ML88], pages 51–60. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Hiraki:1988:EVP

[HSS88] K. Hiraki, S. Sekiguchi, and T. Shimada. Efficient vector processing on a dataflow supercomputer SIGMA-1. In IEEE [IEE88], pages 374–381. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Tang:1988:PVS

[hTDT88] Ju ho Tang, E. S. Davidson, and J. Tong. Polycyclic vector scheduling vs. chaining on 1-port vector supercomputers. In IEEE [IEE88], pages 122–129. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Hatayama:1988:DJS

[HTSH88] S. Hatayama, H. Tsuchiya, Y. Shinkai, and H. Horishige. Development of job-job step scheduler for NAL numerical simulator. In IEEE [IEE88], pages 200– 206. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

IEEE:1988:PSN

[IEE88] IEEE, editor. Proceedings, Supercomputing ’88: November 14– 18, 1988, Orlando, Florida, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v.

2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Inouye:1988:FMD

[IH88] Y. Inouye and S. Hatayama. FA-COM 6443 magneto-optic disk sub-system. In IEEE [IEE88], pages 266–271. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Irani:1988:EBD

[IL88] K. B. Irani and K.-Q. Luc. Elimination of bottlenecks in dynamic dataflow processors. In IEEE [IEE88], pages 80–87. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Ishiguro:1988:VPP

[IMS88] M. Ishiguro, M. Makino, and N. Shinozawa. Vector and parallel processing of the nuclear reactor transient analysis code RELAP5. In IEEE [IEE88], pages 230–236. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Johnston:1988:DSV

[JHH^{+}88] W. E. Johnston, D. E. Hall, J. Huang, M. Rible, Robertson, and D. Distributed scientific video movie making. In IEEE [IEE88], pages 156–162. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Kanada:1988:VMA

[Kan88] Y. Kanada. Vectorization of multiple-precision arithmetic program and 201,326,000 decimal digits of pi calculation. In Martin and Lundstrom [ML88], pages 117–128. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988.

Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Kopetzky:1988:HSH

[Kop88] D. J. Kopetzky. HORSE: a simulation of the Horizon supercomputer. In IEEE [IEE88], pages 53–54. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Kuehn:1988:HSS

[KS88] J. T. Kuehn and B. J. Smith. The Horizon supercomputing system: architecture and software. In IEEE [IEE88], pages 28–34. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Liu:1988:OSP

[LJ88] Y. Liu and O. Johnson. Optimal scheduling policies for mixed scalar-vector multiprocessor supercomputers. In IEEE [IEE88], pages 258–264. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

PeizongLee:1988:HCP

[LK88] Peizong Lee and Z. M. Kedem. On high-speed computing with a programmable linear array. In IEEE [IEE88], pages 425–432. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Lander:1988:SPS

[LMT88] E. Lander, J. P. Mesirov, and W. Taylor, IV. Study of protein sequence comparison metrics on the Connection Machine CM-2. In Martin and Lundstrom [ML88], pages 2–12. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Matzner:1988:ICS

[Mat88] R. A. Matzner. The interaction of cosmic strings — a study of research and teaching graphics. In Martin and Lundstrom [ML88], pages 81–83. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Martin:1988:SPN

[ML88] Joanne L. Martin and Stephen F.

Lundstrom, editors. Supercomputing ’88: proceedings, November 14–18, 1988, Orlando, Florida, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Neves:1988:GDP

[Nev88] K. W. Neves. Growing discord: programming philosophy and hardware design. In IEEE [IEE88], pages 18–26. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Nassersharif:1988:HVA

[NPD88] B. Nassersharif, J. S. Peery, and M. D. DeHart. Highly vectorized algorithm for transient simulation of space reactor systems. In IEEE [IEE88], pages 393–399. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Navon:1988:VCM

[NPR88] I. M. Navon, P. K. H. Phua, and M. Ramamurthy. Vectorization of conjugate-gradient methods for large-scale minimization. In IEEE [IEE88], pages 410– 418. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Obayashi:1988:TSU

[OG88] S. Obayashi and S. Gavali. Threedimensional simulation of underexpanded plumes using upwind algorithms. In Martin and Lundstrom [ML88], pages 25–33. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

ONeil:1988:ASI

[OS88] E. J. O’Neil and C. G. Shaefer. The ARGOT strategy III: the BBN Butterfly multiprocessor. In Martin and Lundstrom [ML88], pages 214–227. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Pryor:1988:VMC

[PB88] D. V. Pryor and P. J. Burns.

Vectorized Monte Carlo molecular aerodynamics simulation of the Rayleigh problem. In IEEE [IEE88], pages 384–391. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Phillips:1988:SVW

[Phi88] R. L. Phillips. A scientific visualization workbench. In IEEE [IEE88], pages 148–155. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Pittelli:1988:ATN

[PS88] F. Pittelli and D. Smitley. Analysis of a 3D toroidal network for a shared memory architecture. In IEEE [IEE88], pages 42–47. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Richards:1988:MSS

[RKZ88] J. Richards, T. Kummell, and D. G. Zarlengo. A mass storage system for supercomputers based on Unix. In IEEE [IEE88], pages 279–286. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Roble:1988:GCS

[Rob88] R. G. Roble. Global change:

the solar influence. In Martin and Lundstrom [ML88], pages 240–250. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Riccardi:1988:AIC

[RS88] G. A. Riccardi and P. H. Schow.

Adaptation of the ISODATA clustering algorithm for vector supercomputer execution. In Martin and Lundstrom [ML88], pages 141–150. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Semtner:1988:BOC

[SC88] A. J. Semtner, Jr. and R. M. Chervin. Breakthroughs in ocean and climate modeling made possible by supercomputers of today and tomorrow. In Martin and Lundstrom [ML88], pages 230–239. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Scheurich:1988:DLC

[SD88a] C. Scheurich and M. Dubois. The design of a lockup-free cache for high-performance multiprocessors. In IEEE [IEE88], pages 352–359. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Stampen:1988:MCC

[SD88b] J. O. Stampen and J. R. Davis.

Multi-issue coalitions in the Congress (voting alignments on supercomputer). In Martin and Lundstrom [ML88], pages 152– 166. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Sato:1988:BCM

[SS88] R. K. Sato and P. N. Swarztrauber. Benchmarking the Connection Machine 2. In IEEE [IEE88], pages 304–309. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Sethian:1988:ISV

[SSG88] J. A. Sethian, J. B. Salem, and A. F. Ghoniem. Interactive scientific visualization and parallel display techniques. In IEEE [IEE88], pages 132–139. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Simmons:1988:PCC

[SW88] Margaret L. Simmons and Harvey J. Wasserman. Performance comparison of the Cray-2 and Cray C-MP/416 supercomputers. In IEEE [IEE88], pages 288–295. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Sterling:1988:TSD

[SWC88] T. L. Sterling, D. S. Wills, and E. Y. Chan. Tokenless static data flow using associative templates. In IEEE [IEE88], pages 70–79. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Swisshelm:1988:MTN

[Swi88] J. M. Swisshelm. Multitasking a three-dimensional Navier– Stokes algorithm on the Cray-2. In Martin and Lundstrom [ML88], pages 14–24. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Tanaka:1988:CTF

[TIGU88] Y. Tanaka, K. Iwasawa, S. Gotoo, and Y. Umetani. Compiling techniques for first-order linear recurrences on a vector computer. In IEEE [IEE88], pages 174–181. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Tsuda:1988:VAV

[TK88] T. Tsuda and Y. Kunieda. V-Pascal: an automatic vectorizing compiler for Pascal with no language extensions. In IEEE [IEE88], pages 182–189. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Thistle:1988:PAH

[TS88] M. R. Thistle and B. J. Smith. A processor architecture for Horizon. In IEEE [IEE88], pages 35–41. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Tomboulian:1988:SSI

[TSM88] S. Tomboulian, C. Streett, and M. Macaraeg. Spectral solution of the incompressible Navier–Stokes equations on the Connection Machine 2. In Martin and Lundstrom [ML88], pages 45–50. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

Whiteside:1988:ULS

[WL88] R. A. Whiteside and J. S. Leichter. Using Linda for supercomputing on a local area network. In IEEE [IEE88], pages 192–199. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Worlton:1988:SPT

[Wor88]  J. Worlton.  Some patterns

of technological change in highperformance computers. In IEEE [IEE88], pages 312–320. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Westbrook:1988:CME

[WP88] C. K. Westbrook and W. J.

Pitz. Computer modeling of engine knock chemistry. In Martin and Lundstrom [ML88], pages 95–104. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. IEEE catalog number 88CH2617-9. IEEE Computer Society Order Number 882.

LingtaoWang:1988:IMI

[WW88] Lingtao Wang and Chuan-Lin Wu. I-NET mechanism for issuing multiple instructions. In IEEE [IEE88], pages 88–95. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Yoshida:1988:TAD

[Yos88]  N. Yoshida.  A transforma-

tional approach to the derivation of hardware algorithms from recurrence equations. In IEEE [IEE88], pages 433–440. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.